put the following equation in standard form, identify center and radius and graph. x squared + y squared+ 6x-4y+12=0
Answer by user_dude2008(1862)

x^2 + y^2+ 6x-4y+12=0


x^2+6x+9-9 + y^2-4y+4-4+12=0

(x+3)^2-9 + (y-2)^2-4+12=0


(x+3)^2+(y-2)^2=1

Circle with center (-3,2) and radius 1